I have the answer to my equation, i just don't know how to solve it
My Equation: 6=1-2n+5
I know the answer is N=0
But i can't remember how to solve it, please help

Answer by nyc_function(2741)

Since the equation has a little n and not a capital N, we are solving for little n.
6 = 1 - 2n + 5
6 = 6 - 2n
6 - 6 = -2n
0 = -2n
0/-2 = n
0 = n
The value of little n is 0.
This is called a linear equation.